Define T:R^3-->R^3 by T(x,y,z)=(-x,x-y,3x+2y+z). Check if T satisfies the polynomial (x-1)(x+1)^2.
1
Expert's answer
2020-03-21T12:28:16-0400

First we have to show that ,"T:R^3 \rightarrow R^3"

Define by "T(x,y,z)=(-x,x-y,3x+2y+z)" is a linear operator.

Let "(x,y,z),(x',y',z') \in R^3" and "a,b \in R"

Where "R" is the underlying field.

Now ,"T\{ a(x,y,z)+b(x',y',z') \}=" "T\{ (ax,ay,az)+(bx',by',bz') \}="

"T" "=\{ (ax+bx',ay+by',az+bz') \}"

"=( - ax-bx,ax-ay+bx'-by',3ax+2ay+az+3bx'+2by'+bz')"

"=(-ax,ax-ay,3ax+2ay+az)" +"(-bx',bx'-by',3bx'+2by'+bz')"

"=a(-x,x-y,3x+2y+z)+b(-x',x'-y',3x'+2y'+z')"

"=aT(x,y,z)+bT(x',y',z')"

Hence "T" is a linear transformation.

Now, the matrix representation of "T" with respect to the usual basis

"\{ (1,0,0),(0,1,0),(0,0,1) \}"is the following:

"\begin{pmatrix}
 -1&0&0\\
1&-1&0\\
3&2&1
\end{pmatrix}"

Since it is a Lower triangular matrix ,its eigen values are "-1,-1,\, and \, 1" .

So its characteristics polynomial is "(x+1)^2(x-1)" .

Again by Cayley Hamilton theorem ,

every square matrix satisfies its characteristic polynomial.

Therefore "T" satisfies "(x-1)(x+1)^2" .